OpenStudy (anonymous):

Passes through (1, 2) with slope -5/3 Write an equation in slope-intercept, point-slope, or standard form for the line with the given information. Explain why you chose the form that you used.

OpenStudy (anonymous):

Point slope would be y-2=-5/3(x-1). Now we can convert to slope intercept. y-2=-5/3x+5/3 +2 on both sides y=-5/3x + 2 5/3 \[-\frac{ 5 }{ 3 }x + 2 \frac{ 5 }{ 3 }\] Now convert from slope intercept to standard form. y=-5/3 + 2 5/3 I would multiply by 3 to clear out the denominator of the x. 3y=-5x+11 Add 5x over 5x+3y=11

OpenStudy (anonymous):

Wow, thank you so much!

OpenStudy (anonymous):

Not a problem. :)